60. Smith Company is attempting to develop the cost function for repair costs. The following past data are available: Machine Hours 4,800 3,400 4,000 5,900 Repair Costs $6,385 4,585 5,285 7,085 Using the high-low method, what is the variable repair cost per machine hour? a) $0.15 b) $1.00 c) $4.00 d) $5.00 Answer: b Difficulty: Easy Learning Objective: Describe cost estimation techniques CPA: Management Accounting/Problem-Solving and Decision-Making Bloomcode: Application 61. Smith Company is attempting to develop the cost function for repair costs. The following past data are available: Machine Hours Repair Costs 4,800 3,400 ,000 ,900 $6,385 4,585 5,285 7,085 Using the high-low method, what is the ixed repair coet? a) $1,185 b) $850 c)$475 d) $565 Answer:a Difficulty: Easy Learning Objective: Describe cost estimation techniques. CPA: Management Aocounting/Problem-Solving and Decision-Making Bloomcode: Application 62. Smith Company is attempting to develop the cost function for repair costs. The following past data are available:
4,800 3,400 4,000 ,900 $6,385 4,585 5,285 7,085 Using the high-low method, what is the estimated repair cost for 4,500 machine hours? a) $5,785 b) $5,585 c) $5,685 d) $5,985 Answer: c Difficulty: Medium Learning Objective: Describe cost estimation techniques. CPA: Management Accounting/Problem-Solving and Decision-Making Bloomcode: Application 63. Milano Company has an average overhead oost per hour of $10.50 at 3,500 machine hours and at 3,000 hours it is $11.25. The company managers wish to estimate the function. What is the variable overhead cost per machine hour? overhead cost a) $1.00 b) $2.00 c) $6.00 d) $8.00 Answer: Difficulty: Easy Learning Objective: Desaribe cost estimation techniques. CPA: Management Aocounting/Problem-Solving and Decision-Making Bloomcode: Applicatiorn 64. Milano Company has an average overhead cost per hour of $10.50 at 3,500 machine hours and at 3,000 hours it is $11.25. The company managers wish to estimate the overhead cost function. What is the fixed overhead cost? a) $15,750 b) $36,750 c) $21,000 d) $18,000 Answer: Difficulty: Easy Learning Objective: Describe cost estimation techniques. CPA: Management Accounting/Problem-Solving and Decision-Making
Answer #1

Answer 60:

Correct answer is:

b.$1.00

Explanation:

Using high low method:

Mchine hour Repair cost High Low Difference 5,900 3,400 2,500 $7,085 $4,585 $2,500

Variable cost per machine hour = $2,500 /2,500 = $1

As such option b is correct and other options a, c and d are incorrect.

Answer 61:

Correct answer is:

a. $1,185

Explanation:

Using high low method:

Mchine hour Repair cost High Low Difference 5,900 3,400 2,500 $7,085 $4,585 $2,500

Variable cost per machine hour = $2,500 /2,500 = $1

Fixed cost = Total repair cost - variable cost = $4,585 - 3,400 * $1 = $1,185

As such option a is correct and other options b, c and d are incorrect.

Answer 62:

Correct answer is:

c. $5,685

Explanation:

As calculated in answer 61 and 62 above:

Fixed cost = $1,185

Variable cost per machine hour = $1

Hence repair cost for 4,500 machine hours =Fixed cost + Variable cost = $1,185 +4,500 * $1 = $5,685

As such option c is correct and other options a, b and d are incorrect.

Answer 63:

Correct answer is:

c. $6

Explanation:

Machine hour Overhead cost per hour Total overhead cost $36,750.00 $33,750.00 $3,000.00 3,500 3,000 500 $10.50 $11.25 Difference

Variable cost per unit = $3,000 / 500 = $6

As such option c is correct and other options a, b and d are incorrect.

Answer 64:

Correct answer is:

a. $15,750

Explanation:

Machine hour Overhead cost per hour Total overhead cost $36,750.00 $33,750.00 $3,000.00 3,500 3,000 500 $10.50 $11.25 Difference

Fixed overhead cost = Total overhead cost - variable overhead cost = $33,750 - 3,000 * $6 = $15,750

As such option a is correct and other options b, c and d are incorrect.